OpenStudy (anonymous):

Type this equation in standard form. 6y - 4 = 3x @Bea_26 @Briaja_Summers @BrittanyDanielle

OpenStudy (anonymous):

if you are talking about the standard form of a linear equation then it's: y=(1/6)(3x+4) or y=(x/2)+(2/3)